# Equitable Allocation for Mixtures of Goods and Chores

### v2

@@ Abstract
Equitable allocation of indivisible items involves partitioning the items among agents such that everyone derives (almost) equal utility. We consider the approximate notion of equitability up to one item (EQ1) and focus on the settings containing mixtures of items (goods and chores), where an agent may derive positive, negative, or zero utility from an item. We first show that—in stark contrast to the goods-only and chores-only settings—an EQ1 allocation may not exist even for additive $\{-1,1\}$ bivalued instances, and its corresponding decision problem is computationally intractable. We focus on a natural domain of normalized valuations where the value of the entire set of items is constant for all agents. On the algorithmic side, we show that an EQ1 allocation can be computed efficiently for (i) $\{-1,0,1\}$ normalized valuations, (ii) objective but non-normalized valuations, (iii) two agents with type-normalized valuations. Previously, EQX allocations were known to exist only for 2 agents and objective valuations, while the case of subjective valuations remained computationally intractable even with two agents. We make progress by presenting an efficient algorithm that outputs an EQX allocation for $\{-1,1\}$ normalized subjective valuations for any number of agents. We complement our study by providing a comprehensive picture of achieving EQ1 allocations in conjunction with economic efficiency notions such as Pareto optimality and social welfare.
@@ 1 Introduction
The distribution of indivisible items fairly among a set of agents is a prominent problem within the field of economics and computation (Moulin, 2004; Brandt et al., 2016; Brams and Taylor, 1996; Varian, 1974). A host of real-world problems deal with the distribution of items that are considered ‘goods’ (positively valued) by some agents while being ‘chores’ (negatively valued) by others: some students may enjoy taking a rigorous course on Artificial Intelligence, while others may find it daunting. Reviewing a paper may be enjoyable if it is aligned with a reviewer’s expertise. Given that agents may have different subjective valuations over the items, the goal is to find an allocation of the items to the agents that is fair while satisfying some notion of economic efficiency.
@@
While an equitable allocation always exists for divisible items (aka cake-cutting) (Alon, 1987; Brams et al., 2006; Cechlárová et al., 2013), when items are indivisible (can not be fractionally assigned), equitable solutions may not exist: consider one good (or chore) and two agents who like (or dislike) it. This impossibility in the presence of indivisible items has motivated ‘up to one’ style approximations. In particular, equitability up to any item (EQX) and equitability up to one item (EQ1) are relaxations that allow for inequality between two agents to be resolved by removing any item (in the case of EQX) or some item (in the case of EQ1) from one agent’s bundle.
In the goods/chores-only settings, existence and computation of EQ1 allocations have been well-explored; such an allocation always exists and can be computed efficiently (Gourvès et al., 2014; Freeman et al., 2020). The setting of mixtures is relatively underexplored. The aim of this exposition is to address this gap. In contrast to the universal existence of EQ1 allocation for goods/chores only settings, such an allocation may not exist for mixtures of goods and chores, even with two agents and two items (Example 3.1).
The above existential issue can be primarily attributed to the pronounced ‘disparity’ in how agents subjectively evaluate the entire set of items. In Example 3.1, Alice’s value for the entire set is $-2$, while Bob’s value is $2$. This observation motivates the study of normalized valuations where the value of the entire set of items is a constant (positive, negative, or zero) for all agents. Normalization is a common and standard assumption in fair division literature, and has been studied in the context of cake-cutting Brandt et al. (2016), [^1] In fair cake-cutting, the common assumption is that each agent values the entire cake as $1$ (or $-1$ for burnt cakes) price of fairness (Caragiannis et al., 2012; Bhaskar et al., 2023) and online fair division as well (Gkatzelis et al., 2021).
@@ 1.1 Our Contributions
Our work provide a deep dive into the existence and computational boundaries of EQ1 allocations for mixtures of goods and chores with and without economic efficiency notions (e.g. Pareto optimality and utilitarian/egalitarian social welfare). We not only show non-existence under non-normalized valuations, but also establish the hardness of deciding the existence of an EQ1 allocation (Theorem 3.2). This result complements a recent study that shows computing an EQX allocation—a strengthening of EQ1—is strongly NP-hard (Barman et al., 2024).
@@
In Section 4, we show that under symmetric tri-valued normalized valuations, where an item takes values from $\{-1,0,1\}$, an EQ1 allocation always exists and can be computed in polynomial time for any number of agents (Theorem 4.4). We note that all our algorithms for $\{-1,0,1\}$ valuations are scale-free and hence, holds for $\{-w,0,w\}$ valuations as well, where $w\in\mathbb{N}$. Such valuations generalize binary preferences, capture realistic scenarios involving approvals/dis-approvals/neutrality and have been studied in the context of EFX$+$PO Aleksandrov and Walsh (2020). It is also relevant to note that binary preferences capture many real-world domains, and many significant contributions in allocation and voting problems have solely focused on such preferences (Bogomolnaia et al., 2005; Babaioff et al., 2021; Halpern et al., 2020; Barman and Verma, 2022; Darmann and Schauer, 2015).
As a corollary, we obtain the existence and computation of EQX allocation for symmetric bi-valued valuations ($\{-1,1\}$). Note that EQX is known to exist only for $2$ agents with objective valuations and is hard to compute for non-objective normalized valuations, even for two agents (Barman et al., 2024). Towards tractability, our algorithm gives an EQX allocation for any number of agents with symmetric and normalized valuations, not necessarily objective. Our techniques involve multiple careful transfers of items among the agents and maintain a partial EQ1 / EQX allocation at every step.
Furthermore, we show that when valuations are in addition type-normalized (i.e. the sum of chores and the sum of goods are independently constant), an EQ1 allocation can be computed efficiently for two agents (Theorem 4.6).
@@
In Section 5, we show an EQ1 and PO allocation may not exist even with $\{-1,1\}$ normalized valuations. Additionally, we show that deciding whether an EQ1$+$PO allocation exists is intractable even for the special class of type-normalized valuations (Theorem 5.2). Nonetheless, we develop a polynomial-time algorithm for $\{-1,0,1\}$ normalized valuations that computes an EQ1$+$PO allocation, when one exists (Theorem 5.3). We note here that the presence of zeros has been a source of computational challenge for EQ1$+$PO allocations even in the only goods setting (Freeman et al., 2019; Garg and Murhekar, 2024). Our algorithmic technique involves reducing the instance to a partial instance such that every item is non-negatively valued by at least one agent, computing a Nash optimal solution, and invoking Lemma 3.5 to assign any remaining chore.
@@
In Section 6, we develop a pseudo-polynomial time algorithms for finding EQX allocations that maximize utilitarian or egalitarian welfare (Theorem 6.1). Note that maximizing welfare along with approximate equitability is already known to be weakly NP-Hard, even when all the items are goods (Sun et al., 2023). In Appendix D, we provide additional results on the compatibility of EQ, EQ1, and PO, with envy-freeness and its relaxations.
@@
Table 1: The summary of our results for arbitrary (non-normalized), objective (non-normalized), normalized, and type-normalized valuations. The result marked by $\star$ and $\dagger^{\prime}$ follow from Freeman et al. (2019) and Sun et al. (2023) respectively. P, (w) NP-c, and (s) NP-h denote polynomial algorithm, weak NP-completeness, and strong NP-hardness. All EQ1 allocations (including welfare-maximizing) can be computed in pseudo-polynomial time (Theorem 6.1). Deciding the existence of EQ1$+$PO allocations is strongly NP-hard even for type-normalized valuations (Theorem 5.2).
@@ 1.2 Related Work
Equitability as a fairness notion was first studied in the context of divisible items, where it is possible to assign a fraction of an item to any agent (Dubins and Spanier, 1961). It is known that equitable allocations always exist for divisible items (Cechlárová et al., 2013; Chèze, 2017). There is a computational limitation to finding such an allocation (Procaccia and Wang, 2017) but approximate equitable allocations admit efficient algorithms (Cechlárová and Pillárová, 2012a, b). In the setting of indivisible items, Freeman et al. (2019, 2020) studied equitability along with efficiency guarantees for goods/chores only settings. Follow-up work has studied equitability in various contexts: when items lie on a path, Follow-up work has studied equitability in various contexts: when items lie on a path, the trade-off between equitability and welfare (the price of equitability), and its interaction with welfare guarantees (Garg and Murhekar, 2024; Suksompong, 2019; Misra et al., 2021; Caragiannis et al., 2012; Aumann and Dombb, 2015; Sun et al., 2023; Bhaskar et al., 2023). Perhaps closest to our work is that of Barman et al. (2024), who study EQX allocations for mixtures and exhibit an algorithm for two agents and objective additive valuations and hardness for non-objective valuations even with two agents.
Mixtures of goods and chores.
Fairness notions like maximin shares (Kulkarni et al., 2021a, b), envy-freeness Bérczi et al. (2020); Hosseini et al. (2023b, a), and competitive equilibrium (Bogomolnaia et al., 2017) have been studied in the context of mixtures Liu et al. (2024). We note here that the setting of mixtures differs substantially from the goods counterpart. Many of the existing algorithmic techniques from the latter fail to work in the former (Aziz et al., 2022; Bhaskar et al., 2020; Bogomolnaia et al., 2017) and hence, the mixed setting demands independent analysis and scrutiny. For the goods-only case, maximizing Nash welfare gives both fairness (envy-freeness up to one item (EF1) [^2] The definitions are deferred to Appendix D.) and efficiency (PO) guarantees (Caragiannis et al., 2019; Barman et al., 2018a) but due to the presence of negative utilities, Nash loses its guarantees in the mixed setting. In addition, for mixtures, EF1+PO is known to exist only for $2$ agents, and the case of arbitrary agents stands as a major open problem (Aziz et al., 2022).
@@ 2 Preliminaries
A fair division instance ($[n],[m],\mathcal{V}$) consists of $n\in\mathbb{N}$ agents, $m\in\mathbb{N}$ items and valuations $V=\{v_{1},v_{2},\ldots v_{n}\}$ where $v_{i}:2^{[m]}\rightarrow\mathbb{Z}$, that captures the value that an agent $i$ derives from a subset of items. We restrict our attention to additive valuations where for any subset $S\subseteq[m]$, we have $v_{i}(S)=\sum_{o\in S}v_{i}\{o\}$. Also, $\{-1,0,1\}$ valuations denote that each item is either valued at a constant $1$, or $-1$, or zero by every agent. An allocation $A=(A_{1},A_{2},\ldots A_{n})$ is a partition of $m$ items into $n$ bundles, one for each agent. An allocation is said to be complete if no item remains unallocated, otherwise it is a partial allocation. We note here that the focus of this work to compute complete allocations.
@@
An allocation $A$ is said to be equitable (EQ) if all the agents derive equal value from their respective bundles, that is, for every pair of agents $i$ and $j$, we have $v_{i}(A_{i})=v_{j}(A_{j})$. An allocation $A$ is said to be equitable up to one item (EQ1) if for any pair of agents $i$ and $j$ such that $v_{i}(A_{i})<v_{j}(A_{j})$, either there exists some good $g$ in $A_{j}$ such that $v_{i}(A_{i})\geq v_{j}(A_{j}\setminus\{g\})$ or there exists some chore $c$ in $A_{i}$ such that $v_{i}(A_{i}\setminus\{c\})\geq v_{j}(A_{j})$. Further, an allocation $A$ is said to be equitable up to any item (EQX) if for any pair of agents $i$ and $j$ such that $v_{i}(A_{i})<v_{j}(A_{j})$, we have $v_{i}(A_{i})\geq v_{j}(A_{j}\setminus\{g\})$ for all goods $g$ in $A_{j}$, and $v_{i}(A_{i}\setminus\{c\})\geq v_{j}(A_{j})$ for all chores $c$ in $A_{i}$.
@@
An allocation $A$ is said to be Pareto dominated by another allocation $A^{\prime}$ if $v_{i}(A^{\prime}_{i})\geq v_{i}(A_{i})$ for all agents $i$ and there exists at least one agent $j$ such that $v_{j}(A^{\prime}_{j})>v_{j}(A_{j})$. If an allocation is not Pareto dominated by any other allocation, then it is said to be Pareto optimal (PO). The utilitarian welfare of an allocation $A$ is the arithmetic mean of individual values under $A$, denoted by $UW(A)=\frac{1}{n}\sum_{i\in[n]}v_{i}(A_{i})$. The egalitarian welfare is the minimum value among all the individual values, denoted by $EW(A)=\min_{i\in[n]}v_{i}(A_{i})$. We denote by UW and EW those allocations that maximize the utilitarian and egalitarian welfare, respectively.
@@
Let $O$ be the set of all items. We say that an item is an objective good if it is valued non-negatively by all the agents. That is, $o\in O$, such that $v_{i}(o)\geq 0$ for all $i\in[n]$ and denote the set of all objective goods by $O^{+}$. Similarly, all items $o\in O$ such that $v_{i}(o)\leq 0$ are called objective chores, denoted by $O^{-}$. The remaining objects are called subjective items, denoted by $O^{\pm}$. We consider two types of normalizations. An instance is normalized if all agents value the entire bundle at a constant, that is, $v_{i}(O)$ is a constant for every $i\in\mathbb{N}$. An instance is type-normalized if for every agent $i$, all the goods sum up to a constant, say $g$, and all the chores sum up to a constant, say $c$, that is, $\sum_{o:v_{i}(o)\geq 0}v_{i}(o)=g$ and $\sum_{o:v_{i}(o)\leq 0)}v_{i}(o)=c$. Note that a type-normalized instance is necessarily normalized (but the converse may not be true).

### v3

@@ Abstract
Equitable allocation of indivisible items involves partitioning the items among agents such that everyone derives (almost) equal utility. We consider the approximate notion of equitability up to one item (EQ1) and focus on the settings containing mixtures of items (goods and chores), where an agent may derive positive, negative, or zero utility from an item. We first show that—in stark contrast to the goods-only and chores-only settings—an EQ1 allocation may not exist even for additive $\{-1,1\}$ bivalued instances, and its corresponding decision problem is computationally intractable. We focus on a natural domain of normalized valuations where the value of the entire set of items is constant for all agents. On the algorithmic side, we show that an EQ1 allocation can be computed efficiently for (i) symmetric tri-valued ($\{-1,0,1\}$) normalized valuations, (ii) objective but non-normalized valuations, (iii) two agents with type-normalized valuations. Previously, EQX allocations were known to exist only for 2 agents and objective valuations, while the case of subjective valuations remained computationally intractable even with two agents. We make progress by presenting an efficient algorithm that outputs an EQX allocation for symmetric bi-valued ($\{-1,1\}$) normalized subjective valuations for any number of agents. We complement our study by providing a comprehensive picture of achieving EQ1 allocations in conjunction with economic efficiency notions such as Pareto optimality and social welfare.
@@ 1 Introduction
The distribution of indivisible items fairly among a set of agents is a prominent problem within the field of economics and computation (Moulin, 2004; Brandt et al., 2016; Brams and Taylor, 1996; Varian, 1974). A host of real-world problems deal with the distribution of items that are considered ‘goods’ (positively valued) by some agents while ‘chores’ (negatively valued) by others: some students may enjoy taking a rigorous course on Artificial Intelligence (AI), while others may find it daunting. Paper reviewing is often rewarding when it aligns with the reviewer’s expertise, but can become a tedious task (a chore) otherwise. Given that agents may have different subjective valuations over the items, the goal is to find an allocation of the items to the agents that is fair while satisfying some notion of economic efficiency.
The integration of modern AI models (e.g. pre-trained Large Language Models) into large-scale decision-making pipelines has become increasingly prevalent, with applications spanning both socio-economic domains and logistical optimization problems. Examples range from recommender systems in streaming platforms to matching mechanisms for course allocation, such as ScheduleScout, [^1] ScheduleScout: https://www.getschedulescout.com/ a platform that has facilitated the allocation of over 400,000 course seats while incorporating satisfaction and fairness criteria. As such systems increasingly mediate the allocation of scarce resources, ensuring equitable outcomes and embedding fairness considerations into their design becomes a critical objective.
The underlying preferences in the above examples can be either positive or negative. The recommender systems, for revenue optimization, may sometimes suggest options that do not align with a user’s preferences. A student may have to be allocated a course that he considers a challenge in order to satisfy the capacity constraints. Therefore, to ensure that no individual/group is systematically disadvantaged in opportunities and outcomes, training equitable and ethical models becomes imperative.
@@
While an equitable allocation always exists for divisible items (aka cake-cutting) (Alon, 1987; Brams et al., 2006; Cechlárová et al., 2013), when items are indivisible (can not be fractionally assigned), equitable solutions may not exist: consider one good (or chore) and two agents who like (or dislike) it. This impossibility in the presence of indivisible items has motivated ‘up to one’ style approximations. In particular, equitability up to any item (EQX) and equitability up to one item (EQ1) are relaxations that allow for inequality between two agents to be resolved by removing any item (in the case of EQX) or some item (in the case of EQ1) from one agent’s bundle (see Section 2 for formal definitions).
In the goods/chores-only settings, existence and computation of EQX/EQ1 allocations have been well-explored; such an allocation always exists and can be computed efficiently (Gourvès et al., 2014; Freeman et al., 2020). The setting of mixtures is relatively underexplored. The aim of this exposition is to address this gap. In contrast to the universal existence of EQX allocation for goods/chores only settings, even the weaker concept of EQ1 allocation may not exist for mixtures of goods and chores, even with two agents and two items, as illustrated in the following example.
Example 1.1.
Consider an instance with two agents (Alice and Bob) and two items such that Alice values the two items at 1-1 each while Bob values them at $1$ each. In any allocation, Alice’s valuation is strictly less than that of Bob even after the hypothetical removal of some item from any of the bundles. Therefore, no allocation in this instance is EQ1.
The above existential issue can be primarily attributed to the pronounced ‘disparity’ in how agents subjectively evaluate the entire set of items. In Example 1.1, Alice’s value for the entire set is 2-2, while Bob’s value is $2$. This observation motivates the study of normalized valuations where the value of the entire set of items is a constant (positive, negative, or zero) for all agents. Normalization is a standard and reasonable assumption in fair division literature, and has been studied in the context of cake-cutting [^2] In fair cake-cutting, the common assumption is that each agent values the entire cake as $1$ (or 1-1 for burnt cakes)(Brandt et al., 2016), price of fairness (Caragiannis et al., 2012; Bhaskar et al., 2023) and online fair division as well (Gkatzelis et al., 2021).
@@ 1.1 Our Contributions
Our work provides a deep dive into the existence and computational boundaries of EQ1 allocations for mixtures of goods and chores with and without economic efficiency notions (e.g. Pareto optimality (PO) and utilitarian/egalitarian social welfare; see Section 2 for formal definitions). We not only show non-existence under non-normalized valuations, but also establish the hardness of deciding the existence of an EQ1 allocation (Theorem 3.1). This result complements a recent study that shows computing an EQX allocation—a strengthening of EQ1—is strongly NP-hard (Barman et al., 2024).
@@
In Section 4, we show that under symmetric bi-valued normalized valuations, where an item takes values from $\{-1,1\}$, an EQX allocation always exists and can be computed in polynomial time for any number of agents (Theorem 4.3). Note that EQX is known to exist only for $2$ agents with objective valuations and is hard to compute for non-objective normalized valuations, even for two agents (Barman et al., 2024). Towards tractability, our algorithm gives an EQX allocation for any number of agents with symmetric and normalized valuations, not necessarily objective. Our techniques involve multiple careful transfers of items among the agents and maintain a partial EQ1/EQX allocation at every step. We additionally show the existence and efficient computation of EQ1 allocations for symmetric tri-valued valuations ($\{-1,0,1\}$). Such valuations generalize binary preferences, capture realistic scenarios involving approvals/dis-approvals/neutrality, and have been studied in the context of EFX++PO (Aleksandrov and Walsh, 2020). It is also relevant to note that binary preferences capture many real-world domains, and many significant contributions in allocation and voting problems have solely focused on such preferences (Bogomolnaia et al., 2005; Babaioff et al., 2021; Halpern et al., 2020; Barman and Verma, 2022; Darmann and Schauer, 2015). We also note that all our results for $\{-1,0,1\}$ valuations are scale-free and hence, hold for $\{-w,0,w\}$ valuations as well, where $w\in\mathbb{R}$.
Furthermore, we show that when valuations are in addition type-normalized (i.e. for all the agents, the sum of the value of all the goods (or chores) is a constant $g$ (or $c$)), an EQ1 allocation can be computed efficiently for two agents (Theorem 4.6).
@@
We also consider EQ1 in conjunction with the efficiency notion of Pareto optimality (PO). An allocation is said to be Pareto optimal if no reallocation of items can make any agent better off without making some other agent worse off. In Section 5, we show that an EQ1 and PO allocation may not exist even with $\{-1,1\}$ normalized valuations. Additionally, we show that deciding whether an EQ1 and PO allocation exists is intractable even for the special class of type-normalized valuations (Theorem 5.2). Nonetheless, we develop a polynomial-time algorithm for $\{-1,0,1\}$ normalized valuations that computes an EQ1++PO allocation, when one exists (Theorem 5.3). We note here that the presence of zeros has been a source of computational challenge for EQ1++PO allocations even in the only goods setting (Freeman et al., 2019a; Garg and Murhekar, 2024). Our algorithmic technique involves reducing the instance to a partial instance such that every item is non-negatively valued by at least one agent, computing a Nash optimal solution, and invoking Lemma 3.4 to assign any remaining chore.
@@
In Section 6, we develop a pseudo-polynomial time algorithm for finding EQX allocations that maximize utilitarian or egalitarian welfare (see Section 2 for formal definitions.) To that end, we use a dynamic programming approach that keeps track of a set of states representing a set of all possible allocations (Theorem 6.1). Note that maximizing welfare along with approximate equitability is already known to be weakly NP-Hard, even when all the items are goods (Sun et al., 2023). In Appendix D, we provide additional results on the compatibility of EQ, EQ1, and PO, with envy-freeness and its relaxations.
@@
Table 1: The summary of our results for arbitrary (non-normalized), objective (non-normalized), normalized, and type-normalized valuations. The result marked by $\star$ follow from (Freeman et al., 2019a) respectively. P, (w) NP-c, and (s) NP-h denote polynomial algorithm, weak NP-completeness, and strong NP-hardness. All EQ1 allocations (including welfare-maximizing) can be computed in pseudo-polynomial time (Theorem 6.1). Deciding the existence of EQ1++PO allocations is strongly NP-hard even for type-normalized valuations (Theorem 5.2).
@@ 1.2 Related Work
Equitability as a fairness notion was first studied in the context of divisible items, where it is possible to assign a fraction of an item to any agent (Dubins and Spanier, 1961). It is known that equitable allocations always exist for divisible items (Cechlárová et al., 2013; Chèze, 2017). There is a computational limitation to finding such an allocation (Procaccia and Wang, 2017), but approximate equitable allocations admit efficient algorithms (Cechlárová and Pillárová, 2012a, b). In the setting of indivisible items, Freeman et al. (2019a, 2020) studied equitability along with efficiency guarantees for goods/chores only settings. Follow-up work has studied equitability in various contexts: when items lie on a path, Follow-up work has studied equitability in various contexts: when items lie on a path, the trade-off between equitability and welfare (the price of equitability), and its interaction with welfare guarantees (Garg and Murhekar, 2024; Suksompong, 2019; Misra et al., 2021; Caragiannis et al., 2012; Aumann and Dombb, 2015; Sun et al., 2023; Bhaskar et al., 2023).
Equitability under non-monotone valuations.
Perhaps closest to our work is that of Barman et al. (2024), who study EQX allocations for mixtures and exhibit an algorithm for two agents and objective additive valuations and hardness for non-objective valuations even with two agents.
Subsequent to our work, Bilò et al. (2025) also noted that EQ1 allocations exist for objective valuations. However, that is the only point of overlap with our work. They studied weaker variants of EQX* and EQ1* in the context of non-monotone valuations. While EQX requires the fairness property to hold regardless of which item is removed, EQX* allows the property to hold for the removal of either goods or chores. EQ1* entails removal of one good from the richer bundle and one chore from the poor bundle. They showed that both EQX* and EQ1* allocations always exists and can be computed in polynomial time.
A similar notion to equitability, named jealousy-freeness (JF), and its up-to-one-style-relaxations JFX and JF1 have also been considered, wherein, in addition to requiring that the hypothetical removals are restricted to non-zero valued items, these definitions also allow hypothetical addition of an item to a bundle (Aleksandrov, 2020). Aleksandrov (2023) segregated the notions where addition/duplication is allowed and renamed these notions to DJFX and DJF1 (that is, JFX and JF1 up to removal or duplication). Their JFX definition is analogous to our EQX definition, albeit they restrict the removals to non-zero valued items. However, we note that their Theorem 3 does not hold true (See Appendix A for details).
Other Fairness Notions.
Fairness notions like envy-freeness (every agent values its own bundle more than it values anyone else’s) (Bérczi et al., 2020; Hosseini et al., 2023b, a) and maximin shares (a threshold based criterion where every agent gets the minimum value it can guarantee itself after dividing the set of items into number of agents many bundles and choosing the worst of them) (Kulkarni et al., 2021a, b) have also been studied in the context of mixtures (Liu et al., 2024). A brief overview of some of the known results in these settings is as follows.
For goods and even monotone valuations, envy-free upto one item (EF1) [^3] See Appendix E for formal definition, which is analogous to EQ1. allocations are always attainable and can be found efficiently via a graph-theoretic envy-cycle elimination algorithm of Lipton et al. (2004). In the chores setting, the standard envy-cycle elimination fails; Bhaskar et al. (2020) overcame this by resolving a top-trading envy cycle, ensuring EF1. For mixed manna, Aziz et al. (2022) established EF1 existence and efficient computation using a double round-robin method.
EF1 is also known to be compatible with PO (Caragiannis et al., 2019) and also admits a pseudo-polynomial time algorithm (Barman et al., 2018a) for goods. For chores, EF1+PO was earlier known to exist only for two agents (Aziz et al., 2022), three agents (Garg et al., 2023), and three agent-types (Garg et al., 2024); and other restricted classes of valuations (Ebadian et al., 2022; Garg et al., 2022; Wu et al., 2023; Garg et al., 2025; Aziz et al., 2023b). The case of four or more agents is very recently settled by Mahara (2025). For mixtures, the joint existence of EF1 and PO allocations remains unresolved even for three agents with additive valuations, except some restricted domains like lexicographic preferences (Hosseini et al., 2023a), identical valuations, and ternary ($\{-\alpha,0,\beta\}$) valuations (Aleksandrov and Walsh, 2020). For two agents, EF1+PO can be achieved via a discrete adjusted winner rule (Aziz et al., 2022). Recently, Barman et al. (2025) showed that envy-freeness up to $k$ reallocations (EFR-$k$) is compatible with PO for any number of agents.
The Maximin Share (MMS) value of an agent is the value that it can guarantee itself if it partitions the set of items into bundles equal to the number of agents, given that it is the last agent to choose its bundle. An allocation that guarantees every agent its MMS value is said to be an MMS allocation. The problem is well-studied when there are only goods (Garg and Taki, 2020; Kurokawa et al., 2018) or chores (Barman and Krishnamurthy, 2020; Huang and Lu, 2021). Notably, an MMS allocation may not even exist for goods with additive valuations (Kurokawa et al., 2018; Feige et al., 2021), but a Polynomial Time Approximation Scheme (PTAS) to compute the MMS values of agents is well-known (Woeginger, 1997). In contrast, for the mixed manna, Kulkarni et al. (2021a) showed that finding even an approximate MMS value of an agent up to any approximation factor in (0,1] is NP-hard. Complementing the hardness, Kulkarni et al. (2021b) gave a PTAS to compute the MMS value, when its absolute value is at least $\frac{1}{p}$ times either the total value of all the goods or total cost of all the chores, for some constant $p\geq 1$. Approximate MMS with PO is also studied, and the problem of computing an $\alpha$-MMS+PO allocation is known to be tractable under the following two conditions: (1) the number of agents is a constant, and (2) for every agent, its absolute value for all the items is at least a constant factor of its total (absolute) value for all the goods or all the chores (Kulkarni et al., 2021a).
@@ 2 Preliminaries
A fair division instance ($n],[m],\mathcal{V}$) consists of $n\in\mathbb{N}$ agents, $m\in\mathbb{N}$ items and valuations $V=\{v_{1},v_{2},\ldots v_{n}\}$ where $v_{i}:2^{[m]}\rightarrow\mathbb{Z}$ captures the value that an agent $i$ derives from a subset of items. We restrict our attention to additive valuations where for any subset $S\subseteq[m]$, we have $v_{i}(S)=\sum_{o\in S}v_{i}\{o\}$. Let $O$ be the set of all $m$ items. We say that an item is an objective good if it is valued non-negatively by all the agents. That is, $o\in O$, such that $v_{i}(o)\geq 0$ for all $i\in[n]$, and denote the set of all objective goods by $O^{+}$. Similarly, all items $o\in O$ such that $v_{i}(o)\leq 0$ are called objective chores, denoted by $O^{-}$. The remaining objects are called subjective items, denoted by $O^{\pm}$. An instance where every item is either an objective good or an objective chore is said to be an instance with objective valuations, otherwise, it is said to be an instance with subjective valuations.
The valuations $\{-1,0,1\}$ denote that each item is either valued at a constant $1$, 0, or 1-1 by every agent and are said to be symmetric tri-valued valuations. Likewise, $\{-1,1\}$ are called symmetric bi-valued valuations. It is relevant to note that all our results for symmetric bi-valued/tri-valued valuations are scale-free and hence, hold for $\{-w,w\}$ and $\{-w,0,w\}$ valuations as well, where $w\in\mathbb{R}$.
An allocation $A=(A_{1},A_{2},\ldots A_{n})$ is a partition of $m$ items into $n$ bundles, one for each agent. An allocation is said to be complete if no item remains unallocated, otherwise it is a partial allocation. The focus of this work is to compute complete allocations.
@@
An allocation $A$ is said to be equitable (EQ) if all the agents derive equal value from their respective bundles, that is, for every pair of agents $i$ and $j$, we have $v_{i}(A_{i})=v_{j}(A_{j})$. An allocation $A$ is said to be equitable up to one item (EQ1) if for any pair of agents $i$ and $j$ such that $v_{i}(A_{i})<v_{j}(A_{j})$, either there exists some good $g$ in $A_{j}$ (a non-negatively valued item by $j$) such that $v_{i}(A_{i})\geq v_{j}(A_{j}\setminus\{g\})$ or there exists some chore $c$ in $A_{i}$ (a non-positively valued item by $i$) such that $v_{i}(A_{i}\setminus\{c\})\geq v_{j}(A_{j})$. Further, an allocation $A$ is said to be equitable up to any item (EQX) if for any pair of agents $i$ and $j$ such that $v_{i}(A_{i})<v_{j}(A_{j})$, we have $v_{i}(A_{i})\geq v_{j}(A_{j}\setminus\{g\})$ for all goods $g$ in $A_{j}$ (non-negatively valued items by $j$), and $v_{i}(A_{i}\setminus\{c\})\geq v_{j}(A_{j})$ for all chores $c$ in $A_{i}$ (non-positively valued items by $i$).
@@
An allocation $A$ is said to be Pareto dominated by another allocation $A^{\prime}$ if $v_{i}(A^{\prime}_{i})\geq v_{i}(A_{i})$ for all agents $i$ and there exists at least one agent $j$ such that $v_{j}(A^{\prime}_{j})>v_{j}(A_{j})$. If an allocation is not Pareto dominated by any other allocation, then it is said to be Pareto optimal (PO). The utilitarian welfare of an allocation $A$ is the arithmetic mean of individual values under $A$, denoted by $UW(A):=\frac{1}{n}\sum_{i\in[n]}v_{i}(A_{i})$. The egalitarian welfare is the minimum value among all the individual values, denoted by $EW(A):=\min_{i\in[n]}v_{i}(A_{i})$. The Nash welfare is the geometric mean of the individual values, denoted by $NW(A):=(\prod_{i\in[n]}v_{i}(A_{i}))^{\frac{1}{n}}$ We denote by UW, EW and NW those allocations that maximize the utilitarian, egalitarian, and Nash welfare, respectively.
@@
We consider two types of normalizations. An instance is normalized if all agents value the entire bundle at a constant, that is, $v_{i}(O)$ is a constant for every agent $i\in\mathbb{N}$. An instance is type-normalized if for every agent $i$, all the goods sum up to a constant, say $g$, and all the chores sum up to a constant, say $c$, that is, $\sum_{o:v_{i}(o)\geq 0}v_{i}(o)=g$ and $\sum_{o:v_{i}(o)\leq 0}v_{i}(o)=c$. Note that a type-normalized instance is necessarily normalized (but the converse may not be true).